OpenStudy (anonymous):

5(2+4z) = 85 (PLEASE SHOW WORK, THANKS)

OpenStudy (anonymous):

5(2+4z)=85 (5/5)(2+4z)=85/5 2+4z=17 4z-2+2=17-2 4z=15 z=15/4

OpenStudy (anonymous):

2 + 4z = 17 4z = 15 z = 15/4

OpenStudy (anonymous):

THANK YOU BOTH!!!! SOOOOOOOOOOOOOOOO MUCH!!!